Order: Dobutrex 3 mcg/kg/min Available: Dobutrex 500 mg in 250 mL D5 1/2 NS Clients weight: 149.6 lb The IV flow rate is
Answer #1

Ans) Client's weight = 149.6 lbs = 67.857 kg ( 1 lbs /2.205 = kg)

3mcg = 0.003 mg ( 1 mcg/1000 = 1 mg)

0.003× 67.857 = 0.2 mg/min

500 mg = 250 ml

0.2 mg = x

x = 0.2× 250/ 500 = 0.1ml/min

The IV flow rate to deliver this dosage is 0.1 ml/min.